OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32 A. an = -4 + an-1 B. an = -2 + an-1 C. an = -2 • an-1 D. an = -4 • an-1

OpenStudy (amistre64):

geometric have nothing to do with addition or subtraction

OpenStudy (amistre64):

the general formula for a geometric seq is:\[a_n=r*a_{n-1}\]

OpenStudy (amistre64):

so to figure out r; just find the value which get the same outcome an = r * an-1 8 = r * -2